# QuNu Collaboration

### v2

@@ I Introduction
Spontaneous chiral symmetry breaking is a fundamental phenomenon in Quantum Chromodynamics (QCD), which governs the strong interactions between quarks and gluons [1]. The chiral condensate, formed by the quark-antiquark pairs due to spontaneous chiral symmetry breaking, imparts a non-zero expectation value to the bilinear fermion operator $\bar{\psi}\psi$, which is crucial for comprehending the origin of quark mass [2, 3, 4, 5, 6, 7]. In QCD systems with high temperature, quarks and gluons deconfine, leading to the formation of quark-gluon plasma (QGP) and chiral symmetry restoration [8]. On the other hand, under high density conditions, quarks are highly overlapping, and the ordinary quark-antiquark condensates are replaced by color-flavor locking of quark Cooper pairs [9, 10]. The restoration of chiral symmetry at high temperature and density is crucial to understanding the evolution of the early universe, as well as the properties of QGP and neutron stars [11, 12, 13, 14].
Due to the inherently non-perturbative nature of QCD, the first principle theoretical and numerical study of chirality imbalance present significant challenges [15, 16]. For finite temperature systems, lattice gauge theory (LGT) has emerged as a prominent theory in addressing these challenges[17, 18]. In LGT, as constrained by gauge invariance to facilitate quark-gluon interactions, the matter degrees of freedom are encoded into the lattice sites and the gauge degrees of freedom are represented by the links between different lattice sites. Monte Carlo methods, employed in LGT, provide a direct approach for solving non-perturbative QCD problems from first principles and have achieved significant success [14, 19, 20]. However, such a method encounters the notorious sign problem [21] when addressing issues related to finite chemical potential for high density systems. Though Taylor expansion techniques have been performed to alleviate the sign problem for small chemical potential systems [22], the issue in high chemical potential systems remains unresolved.
In recent years, with the rapid development of quantum hardware, quantum computing has become increasingly prominent in dealing with the sign problems in high-energy physics [23, 24, 25], such as nucleon structure [26, 27, 28, 29] and parton fragmentation [30, 31]. In addition, another major advantages of quantum simulation is that it enables the study of thermal equilibrium systems while circumventing the sign problem associated with traditional Monte Carlo methods [32, 33, 34, 35, 36, 37, 38]. In particular, quantum simulations for real-time chiral dynamics at finite temperature have been investigated in 1+1 dimensional Nambu-Jona-Lasinio model [32] and Schwinger model [39] using quantum imaginary time evolution algorithm.
@@
We simulate explicitly the chiral condensate in finite temperature and density through the evaluation of order parameter for spontaneous chiral symmetry breaking, using the 1+1 dimensional SU(2) non-Abelian gauge theory. To obtain the thermal average of the chiral condensate, it is essential to efficiently prepare the Gibbs state on a quantum computer. Given that the Gibbs state corresponds to the minimum of the free energy, the strategy involves preparing the Gibbs state is minimizing the free energy by variational methods. The Quantum Alternating Operator Ansatz (QAOA) [45], known for its robust representation power, is utilized to parameterize the free energy. We observe spontaneous chiral symmetry breaking at low temperatures and automatic restoration at high temperatures within the non-Abelian theoretical framework by variational quantum algorithm, which is consistent with theoretical expectations. We also investigate the behavior of chiral condensate at different chemical potentials. To validate the effectiveness of our variational method, we compare our simulation results with those obtained from exact diagonalization. Our work demonstrates the potential of using near-term quantum computers to explore QCD systems with finite temperature and finite density.
@@ II.1 SU(2) non-Abelian gauge theory on the lattice
For the purposes of facilitating quantum simulation, we discretize the continuous version of the Hamiltonian using staggered fermions. The Kogut-Susskind formulation of the Yang-Mills Hamiltonian[18, 47] is given by
@@
The last term of the Hamiltonian represents the energy of the chromoelectric field, in which the chromoelectric field $\mathbf{L}_{n}$ can be expressed through its left chromoelectric field or right chromoelectric field components, $\mathbf{L}_{n}^{2}=L_{n}^{a}L_{n}^{a}=R_{n}^{a}R_{n}^{a}$. The left and right chromoelectric fields are related through the adjoint representation of the SU(2) group, $R_{n}^{a}=\left(U_{n}^{{\rm adj}}\right)^{ab}L_{n}^{b}$ [^2] The matrix element of the SU(2) group in the adjoint representation is given by: $(U_{n}^{{\rm adj}})^{ab}=2Tr[U_{n}t^{a}U_{n}^{\dagger}t^{b}]$, and satisfy the algebra
@@ II.2 Purely fermionic formulation and qubit encoding
Before proceeding, for clarity in representation, we rewrite the Hamiltonian was rewritten
@@
$H^{\prime}\rightarrow\Theta H\Theta^{\dagger}$ $=\frac{1}{2}\sum_{n=0}^{N-2}\left(\phi_{n}^{\dagger}\phi_{n+1}+H.C.\right)$ (11)
@@
$\langle\bar{\psi}\psi\rangle=\sum_{n=0}^{N-1}\left[\left(-1\right)^{n+1}(\sigma_{2n}^{+}\sigma_{2n}^{-}+\sigma_{2n+1}^{+}\sigma_{2n+1}^{-})\right].$ (18)
@@ II.3 Variational quantum algorithm for chirality imbalance
where $\rho$ is the density matrix and $Z$ is the partition function with $\beta$ the inverse of temperature. Since the Gibbs state corresponds to the minimum of the system’s free energy, our motivation to obtain the Gibbs state is to minimize the free energy through variational methods. The free energy of a system at finite temperature can be expressed as
@@
The density matrix of Gibbs state has a simple form when expressed by the energy eigenstates $|i\rangle$
@@
where each state’s probability is determined by the corresponding energy $P_{i}=e^{-\beta E_{i}}/\sum_{n}e^{-\beta E_{n}}$. In this expression, only the $P_{i}$s depend on the temperature, and once the eigenstates are obtained, $P_{i}$s can be calculated easily. Therefore, we propose a parametrized mixed state as
@@
where $|\varphi_{i}\rangle$ represents an initial pure state that can be readily prepared on a quantum computer. The probability $P_{i}$ is then given by $P_{i}=\exp\left[-\beta\langle\varphi_{i}|U(\theta)^{\dagger}HU(\theta)|\varphi_{i}\rangle\right]/Z$. The parameters are only related to the states, which means they do not depend on the temperature. Therefore, once they are optimized at one temperature, they are fixed throughout the following calculation.
@@
$U(\theta)=\prod_{i=1}^{p}\prod_{j=1}^{n}\exp\left(i\theta_{ij}H_{j}\right),$ (23)
where the Hamiltonian is decomposed into $H=H_{1}+H_{2}+\cdots+H_{n}$, with $n\geq 2$. Each component $H_{i}$ preserves the same symmetries as the full Hamiltonian $H$ and satisfies the commutation relation $\left[H_{i},H_{j}\right]\neq 0$ with every other component. $p$ is the number of layers in the quantum circuit. Generally, the greater the values of $p$ and $n$, the stronger the representation power of the ansatz.
@@
    Randomly select a set of states $|\varphi_{i}\rangle$, and construct the initial mixed state from the chosen pure states. Then minimize the free energy with regard to the parameterized mixed state, and output the final parameters $\mathbf{\theta}$.
@@
    Update the mixed state by choosing another set of states. Using the previously obtained parameters as initial parameters to perform a variation on the free energy, and subsequently obtain updated parameters.
@@
We have tested that the required number of configurations in each step above is not extensive. Taking 8 qubits as an example, a set of 20 states for each step suffices to represent the density matrix of the Gibbs state.
After obtaining the optimized $U(\theta)$, we can derive any eigenstate and its eigenvalue of the system. In particular, we perform a Monte Carlo sampling to get the desired density matrix. The steps for Monte Carlo sampling are outlined below
@@
    Initialize a state in computational basis $|\varphi_{i}\rangle$, calculate the eigenstate $U(\theta)|\varphi_{i}\rangle$ and its energy $E_{i}$. In practice, we always start from the ground state $|\varphi_{0}\rangle$.
@@
The above process yields a set of states that reflect the thermodynamic properties of the system, which allows us to calculate the thermal average of an observable $O$
@@ III Results
Figure 1: The chiral condensate as a function of temperature $T/g$, in the case of $m/g\in\{5,10,15,20,25,30\}$ with a system size of 8 qubits, where all configurations are considered.
@@
Figure 2: The chiral condensate as a function of temperature $T/g$, in the case of $m/g\in\{5,10,15,20,25,30\}$, with a system size of 8 qubits, which is obtained from MC calculation.
Figure 3: The chiral condensate as a function of temperature $T/g$, in the case of $m/g=5$ with a system size of 12 qubits. The left figure shows the results of sampling 1,000 states, while the right figure presents the results from sampling 2,000 states.
Figure 4: The chiral condensate as a function of temperature $T/g$, in the case of $m/g=5$, with a system size of 8 qubits, at different $\mu/g\in\{4,6,8,10,20,30,40\}$.
@@
To test our proposed algorithm on real quantum hardware, we execute our quantum algorithm on IBM’s real quantum hardware. Owing to the current limitations of quantum resources, we adopted an alternative ansatz as follows
@@
This ansatz significantly reduces the circuit depth at the cost of no longer preserving quantum numbers during time evolution. Additionally, due to time constraints, the optimization of parameters is still carried out on a classical computer. After the parameters are optimized, we compute the chiral condensate using all eigenstates of the system on IBM’s quantum hardware. The comparison with exact diagonalization is shown in Fig. 5, and a reasonable good agreement between real quantum computation and exact diagonalization is observed. Furthermore, we compare the eigenenergy obtained from quantum hardware and exact diagonalization. The results for the relative error are displayed in Fig. 6. In this figure, computational bases are renumbered from 0 to 255 based on the magnitude of their energies. As we can see, the results obtained from IBM’s quantum hardware are in good agreement with the ones from exact diagonalization. Notice that the large relative errors for low energy states arise from the addition of a constant to the Hamiltonian, leading to the normalization of the strong coupling ground state energy to zero. The agreement between actual quantum hardware and exact diagonalization indicates the feasibility of using near-term quantum computers for the study of chiral condensation.

### v3

@@ I Introduction
Spontaneous chiral symmetry breaking is a fundamental phenomenon in Quantum Chromodynamics (QCD), which governs the strong interactions between quarks and gluons [1]. The chiral condensate, formed by the quark-antiquark pairs due to spontaneous chiral symmetry breaking, imparts a non-zero expectation value to the bilinear fermion operator $\bar{\psi}\psi$, which is crucial for comprehending the origin of quark mass [2, 3, 4, 5, 6, 7]. In QCD systems with high temperature, quarks and gluons deconfine, leading to the formation of quark-gluon plasma (QGP) and chiral symmetry restoration [8]. On the other hand, under high-density conditions, quarks are highly overlapping, and the ordinary quark-antiquark condensates are replaced by color-flavor locking of quark Cooper pairs [9, 10]. The restoration of chiral symmetry at high temperature and density is crucial to understanding the evolution of the early universe, as well as the properties of QGP and neutron stars [11, 12, 13, 14].
Due to the inherently non-perturbative nature of QCD, the first principle theoretical and numerical studies of chirality imbalance present significant challenges [15, 16]. For finite temperature systems, lattice gauge theory (LGT) has emerged as a prominent theory in addressing these challenges[17, 18]. In LGT, as constrained by gauge invariance to facilitate quark-gluon interactions, the matter degrees of freedom are encoded into the lattice sites and the gauge degrees of freedom are represented by the links between different lattice sites. Monte Carlo methods, employed in LGT, provide a direct approach for solving non-perturbative QCD problems from first principles and have achieved significant success [14, 19, 20]. However, such a method encounters the notorious sign problem [21] when addressing issues related to finite chemical potential for high density systems. Though Taylor expansion techniques have been performed to alleviate the sign problem for small chemical potential systems [22], the issue in high chemical potential systems remains unresolved.
In recent years, with the rapid development of quantum hardware, quantum computing has become increasingly prominent in dealing with the sign problems in high-energy physics [23, 24, 25, 26], such as nucleon structure [27, 28, 29, 30] and parton fragmentation [31, 32]. In addition, another major advantages of quantum simulation is that it enables the study of thermal equilibrium systems while circumventing the sign problem associated with traditional Monte Carlo methods [33, 34, 35, 36, 37, 38, 39]. In particular, quantum simulations for real-time chiral dynamics at finite temperature have been investigated in 1+1 dimensional Nambu-Jona-Lasinio model [33], SYK model[40] and Schwinger model [41] using quantum imaginary time evolution algorithm[42].
@@
We simulate explicitly the chiral condensate in finite temperature and density through the evaluation of order parameter for spontaneous chiral symmetry breaking, using the 1+1 dimensional SU(2) non-Abelian gauge theory. To obtain the thermal average of the chiral condensate, it is essential to efficiently prepare the Gibbs state on a quantum computer. Given that the Gibbs state corresponds to the minimum of the free energy, the strategy involves preparing the Gibbs state is minimizing the free energy by variational methods. The Quantum Alternating Operator Ansatz (QAOA) [48], known for its robust representation power, is utilized to parameterize the free energy. The Monte Carlo method is commonly used in lattice field theory to simulate thermal systems effectively. In particular, we implement the Monte Carlo random sampling for the optimization of eigenstates at a specific temperature, which also suffices for any other temperatures, significantly reducing the computational cost. Additionally, the Monte Carlo method is used to calculate the desired density matrix, offering a substantial advantage for large qubit systems. We observe spontaneous chiral symmetry breaking at low temperatures and automatic restoration at high temperatures within the non-Abelian theoretical framework by variational quantum algorithm, which is consistent with theoretical expectations. We also investigate the behavior of chiral condensate at different chemical potentials. To validate the effectiveness of our variational method, we compare our simulation results with those obtained from exact diagonalization. Our work demonstrates the potential of using near-term quantum computers to explore QCD systems with finite temperature and finite density.
@@ II.1 SU(2) non-Abelian gauge theory on the lattice
For the purposes of facilitating quantum simulation, we discretize the continuous version of the Hamiltonian using staggered fermions. The Kogut-Susskind formulation of the Yang-Mills Hamiltonian[18, 50][51, 52], is given by
@@
The last term of the Hamiltonian represents the energy of the chromoelectric field, in which the chromoelectric field $\mathbf{L}_{n}$ can be expressed through its left chromoelectric field or right chromoelectric field components, $\mathbf{L}_{n}^{2}=L_{n}^{a}L_{n}^{a}=R_{n}^{a}R_{n}^{a}$. The left and right chromoelectric fields are related through the adjoint representation of the SU(2) group, $R_{n}^{a}=\left(U_{n}^{{\rm adj}}\right)^{ab}L_{n}^{b}$ [^2] The matrix element of the SU(2) group in the adjoint representation is given by: $(U_{n}^{{\rm adj}})^{ab}=2{\rm Tr}[U_{n}t^{a}U_{n}^{\dagger}t^{b}]$, and satisfy the algebra
@@ II.2 Purely fermionic formulation and qubit encoding
Before proceeding, for clarity in representation, we rewrite the Hamiltonian as
@@
$H^{\prime}\rightarrow\Theta H\Theta^{\dagger}$ $=\frac{1}{2}\sum_{n=0}^{N-2}\left(\phi_{n}^{\dagger}\phi_{n+1}+{\rm H.C.}\right)$ (11)
@@
$\langle\bar{\psi}\psi\rangle=\left\langle\sum_{n=0}^{N-1}\left[\left(-1\right)^{n+1}(\sigma_{2n}^{+}\sigma_{2n}^{-}+\sigma_{2n+1}^{+}\sigma_{2n+1}^{-})\right]\right\rangle.$ (18)
@@ II.3 Variational quantum algorithm for chirality imbalance
where $\rho$ is the density matrix and $Z$ is the partition function with $\beta$ the inverse of temperature. Since the Gibbs state corresponds to the minimum of the system’s free energy, our motivation to obtain the Gibbs state is to minimize the free energy through variational methods [62]. The free energy of a system at finite temperature can be expressed as
@@
However, if one directly parameterizes the density matrix, the parameters will usually depend on the temperature, and therefore, the optimization must be done at each different temperature. Here, we propose a modified algorithm taking advantage of the Gibbs state being diagonal when expressed by the energy eigenstates $|i\rangle$
@@
where each state’s probability is determined by the corresponding energy $P_{i}=e^{-\beta E_{i}}/\sum_{q}e^{-\beta E_{q}}$. In this expression, only the $P_{i}$s depend on the temperature, and once the eigenstates are obtained, $P_{i}$s can be calculated easily. Therefore, we propose a parametrized mixed state as
@@
where $|\varphi_{i}\rangle$ represents initial states from a orthonormal basis. The probability $P_{i}$ is then given by $P_{i}=\exp\left[-\beta\langle\varphi_{i}|U(\theta)^{\dagger}HU(\theta)|\varphi_{i}\rangle\right]/Z$. Since $U(\theta)$ is a unitary operator, the entropy can then be calculated as $S=-\sum_{i}P_{i}\log P_{i}$, which avoids computing the logarithm of the density matrix. The parameters are only related to the energy eigenstates, which means they do not depend on the temperature. Therefore, once they are optimized at one temperature, they are fixed throughout the following calculation.
@@
$U(\theta)=\prod_{i=1}^{p}\prod_{j=1}^{d}\exp\left(i\theta_{ij}H_{j}\right),$ (23)
where the Hamiltonian is decomposed into $H=H_{1}+H_{2}+\cdots+H_{d}$, with $d\geq 2$. Each component $H_{i}$ preserves the same symmetries as the full Hamiltonian $H$ and satisfies the commutation relation $\left[H_{i},H_{j}\right]\neq 0$ with every other component. $p$ is the number of layers in the quantum circuit. Generally, the greater the values of $p$ and $d$, the stronger the representation power of the ansatz.
@@
    Randomly select a set of $M$ states $|\varphi_{i}\rangle$ in an orthonormal basis (here we use the computational basis for simplicity), and construct the initial mixed state from the chosen pure states. Then, minimize the free energy with regard to the parameterized mixed state and output the final parameters $\mathbf{\theta}$.
@@
    Update the mixed state by choosing another set of $M$ states. Using the previously obtained parameters as initial parameters to perform a variation on the free energy, and subsequently obtain updated parameters.
@@
The advantage of the QAOA is that $U(\theta)$ preserves the same symmetries as $H$. This ensures that all states that are related by conserved symmetries are optimized simultaneously, thus accelerating the optimization. We have tested that the required number of configurations in each step above is not extensive. Taking 8 qubits as an example, $M=20$ for each step suffices to represent the density matrix of the Gibbs state, which is more efficient than the calculation considering the full 256 eigenstates.
Increasing the number of states used in each step will improve the accuracy of optimization while increasing the computational cost. In addition, using a small set of states can introduce stochasticity and help to avoid falling into local minima during optimization. This strategy is similar to the mini-batch gradient descent method in machine learning, where a subset of data is used in each iteration to balance computational efficiency and optimization performance [64].
After obtaining the optimized $U(\theta)$, we perform a Monte Carlo sampling to get the desired density matrix. The steps for Monte Carlo sampling are outlined below
@@
    Initialize a state from the same set of orthonormal basis as used in the optimization steps $|\varphi_{i}\rangle$, calculate the eigenstate $U(\theta)|\varphi_{i}\rangle$ and its energy $E_{i}$. In practice, we always start from the state $|\varphi_{0}\rangle$ that is mapped to the ground state.
@@
In principle, since the unitary operator $U$ is optimized by a random process, states that are not selected in the optimization steps are not guaranteed to be optimized. However, as discussed above, the symmetry preserving $U$ helps to extend the set of optimized states. Also, since eigenstates with lower energies contribute more to the density matrix, both the optimization and the sampling algorithm will naturally focus on these states; therefore, we do not restrain the states that can be selected in the sampling.
@@
The above process yields a set of states that effectively reflect the thermodynamic properties of the system, which allows us to calculate the thermal average of an observable $O$
@@ III Results
Figure 1: The chiral condensate as a function of temperature $T/g$, in the case of $ma,m/g\in\{5,10,15,20,25,30\}$ with a system size of 8 qubits, where all configurations are considered.
@@
The statistical errors can be estimated by standard techniques like bootstrapping. However, one must take into consideration that the samples generated by our Monte Carlo method are autocorrelated, so the effective sample size is reduced. Overall, the errors are consistent with the deviation between numerical and theoretical results. Also, on a state-of-the-art quantum computer, the statistical errors should be less important compared to the errors from quantum circuits. Therefore, the statistical errors are not shown in our results.
Figure 2: The chiral condensate as a function of temperature $T/g$, in the case of $ma,m/g\in\{5,10,15,20,25,30\}$, with a system size of 8 qubits, which is obtained from MC calculation using 1000 sampled states.
Figure 3: The chiral condensate as a function of temperature $T/g$, in the case of $ma,m/g=5$ with a system size of 12 qubits. The left figure shows the results of sampling 1,000 states, while the right figure presents the results from sampling 2,000 states.
Figure 4: The chiral condensate as a function of temperature $T/g$, in the case of $ma,m/g=5$, with a system size of 8 qubits, at different $\mu/g\in\{4,6,8,10,20,30,40\}$.
@@
To test our proposed algorithm on real quantum hardware, we execute our quantum algorithm on IBM’s real quantum hardware, specifically the ibm_sherbrooke device. Owing to the current limitations of quantum resources, we adopted an alternative ansatz as follows
@@
This ansatz significantly reduces the circuit depth at the cost of no longer preserving quantum numbers during time evolution. Additionally, due to time constraints, the optimization of parameters is still carried out on a classical computer. After the parameters are optimized, we compute the chiral condensate using all eigenstates of the system on IBM’s quantum hardware, with 5000 shots per measurement. The comparison with exact diagonalization is shown in Fig. 5, and a reasonable good agreement between real quantum computation and exact diagonalization is observed. Furthermore, we compare the eigenenergy obtained from quantum hardware and exact diagonalization. The results for the relative error are displayed in Fig. 6. In this figure, computational bases are renumbered from 0 to 255 based on the magnitude of their energies. As we can see, the results obtained from IBM’s quantum hardware are in good agreement with the ones from exact diagonalization. Notice that the large relative errors for low energy states arise from the addition of a constant to the Hamiltonian, leading to the normalization of the strong coupling ground state energy to zero. The agreement between actual quantum hardware and exact diagonalization indicates the feasibility of using near-term quantum computers for the study of chiral condensation.